OpenStudy (anonymous):

Find the equation of the quadratic function with zeros -4 and 4 and vertex at (0, -16)

OpenStudy (anonymous):

help

OpenStudy (anonymous):

how

OpenStudy (anonymous):

you can plug in either zero into the vertex form

OpenStudy (anonymous):

is the answer y = x^2 - 16 or y = -x^2 - 16

OpenStudy (anonymous):

Vertex form: $$ \Large y = a(x-h)^2 + k $$ y = a ( x - 0)^2 - 16 y = a x^2 - 16 now plug in (4,0) or (-4,0) and solve for `a`

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you should get a = 1

OpenStudy (anonymous):

so the answer is y= x^2 -16 ?